If 17.0 gallons of gasohol contains 1.75 gallons of ethyl alcohol, what is the percent of alcohol in the gasohol?


1
Expert's answer
2019-10-08T07:42:47-0400

The percentage of alcohol in a gasohol by volume is

"\omega = \frac{V(alcohol)}{V(gasohol)} \cdot 100 \% =\frac{1.75 gal}{17.0gal} \cdot 100 \% = 10.3 \ vol\%"